# WaveletGPT: Wavelet Inspired LLMs

### v2

@@ 5 Long Range Arena Benchmarks
We use the setup provided by Khalitov et al. (2022) to extract the data and be uniform with other benchmarks. We use a nearly identical architecture for all three modalities, only modifying the embedding matrix to accommodate different tokenizers and output categories. Our baseline consists of a 6-layer causal Transformer decoder with a model dimension of 32 and a feed-forward dimension four times that of the embedding dimension. We extract the last token of the sequence as a 32-dimensional embedding for classification, followed by a dense layer with 2048 neurons and another dense layer corresponding to the number of categories. The input goes through an embedding layer that converts discrete tokens into a 32-dimensional vector. The input vocabularies are 256 for text and image, and 16 for ListOps. The context lengths are 2048, 3072, and 1999 tokens, respectively, with output categories of 2, 10, and 10. In our modified architecture, we introduce our waveletGPT module between each decoder layer, retaining half of the embedding dimensions as they are. For the other half, we use non-learnable kernels, increasing the kernel size from 2, 4, and 8 to 512 linearly for dimensions 16 to 32, while maintaining the causality assumption. This introduces highways that hierarchically process data at each embedding and Transformer decoder layer without adding parameters, similar to our approach for LLM. As shown in Table 2, we achieve notable gains across all three modalities, where even small improvements are worth reporting. We significantly outperform non-causal methods, such as (Zhuang et al., 2024), with nearly 2% improvement on ListOps and 4.5% on a much smaller architecture—ours has 32 dimensions and six layers compared to 128 dimensions and eight layers. We limit our comparison method for fairness only with vanilla Transformer architectures. We also compare with two non-casual architectures that incorporated signal processing based ideas: FNET and WavSPA. We do not compare it with other sophisticated state space based methods or complex architectural changes as it would have required further tuning to our method/ significant architectural changes than straightforward simple tweaks to have a fair comparison. Compared to non-causal FNet, our model significantly outperformed all three LRA tasks, achieving 20% improvement on ListOps and Image and 10% on text. The most notable gain is in the ListOps task, which involves modeling a hierarchical, tree-like structure of math operations, making our model particularly suitable. To the best of our knowledge and Liu et al. (2024), this is the best performance achieved by a simple attention-based Transformer architecture on LRA tasks.
@@ 6 Conclusion and Future Work
We showcase the powerful incorporation of a core signal processing idea, namely wavelets, into large language model pre-training. By imposing a multi-scale structure onto every intermediate embedding, we achieve the same performance 40-60% faster, compared to a baseline architecture. We achieve a substantial performance boost if we train for the same number of steps. Our method generalizes across three modalities: raw text, symbolic music, and raw audio, giving similar performance speedups. Several exciting directions can be explored in future work, including incorporating more advanced ideas from wavelets and multi-resolution signal processing onto large language models. It will be interesting to see how the model behaves for different variants of multi-scale structures.
7 Acknowledgement
@@
This work was supported by the Stanford Institute of Human-Centered AI (HAI) through a Google Cloud computing grant. The author is thankful to Prof. Dan Jurafsky for feedback and helpful discussions on the problem and its impact. I am also thankful to the anonymous reviewers for various suggestions.
Appendix A Comparison With Exponential Moving Averages
We compare our method with Exponential Moving Averages (EMA) on the intermediate signals. This is widely used in time-series analysis for smoothening data, and it is another type of way that can modify intermediate signals. We proposed Haar wavelet, a multi-resolution kernel that can look at the input signal at various levels of scales depending on embedding dimension. We will now compare it against an EMA baseline and motivate where we differ and are similar to our proposed method.
A.1 Background
Loosely speaking, instead of a moving average filter taking the mean of the signal, an EMA uses a different kernel, i.e., an exponential function. Meanwhile, a moving average kernel assigns equal weight to all time points. If we assume that the $x_{i}^{l}(t)$ signal of length is equal to context length after the $l^{th}$ layer with $t$ being the token index going from 0 to context length $L$ at embedding dimension $i$, we can define the modified exponential smoothed version of the signal $s_{t}$ as
$s_{0}=x_{i}^{l}(0)\quad s_{t}=\alpha x_{i}^{l}(t)+(1-\alpha)s_{t-1}$
Where $\alpha$ is the decay factor, it always satisfies $10<\alpha<$ 1. We can observe that for each of the tokens, depending on the decay factor $\alpha$, we assign weights to the more recent values over the past values. When $\alpha=1$, the weightage given is only to the current observation, and when $\alpha=0$, it is just flat and gives equal weightage. The differences with moving average filters are evident i.e., first, the moving average filter gives equal weight to all of the values in a window to update the values of a particular window. Depending on the value of $\alpha$, an EMA filter gives an exponential weighted kernel. However, from the definition itself, an EMA filter, irrespective of the value of $\alpha$, is an Infinite-Impulse response (IIR) filter, whereas a moving average filter is a finite impulse response (FIR) filter. Therefore, for every value update at a particular location, the values of dependencies of the previous samples will never be zero and relatively small. One can see that these values can add up significantly for some values of $\alpha$ when we are predicting the next tokens at longer context lengths. Due to the nature of the IIR filter, the values are never zero. They are assigned values weighed depending on the previous observation as 1, $1-\alpha$, $(1-\alpha)^{2}$, $(1-\alpha)^{3}$,…
On the other hand, our proposed method includes wavelets composed mainly of FIR filters, including Haar or Daubechies. They are, therefore, only limited to a finite duration and can be adapted in multi-resolution setups with varied window lengths, as we have proposed in our paper. This allows us to have multi-scale information where we look at any signal at different resolutions with varied window lengths, with no contributions from components outside the desired window. (as we set the contribution from those values as 0). EMA, on the other hand, would still have some contribution from every component due to its recursive nature. One could also have a version similar to our method where one could vary $\alpha$ depending on the embedding dimension $i$. The update equations would now be a function of $i$, i.e.
$s_{0}=x_{i}^{l}(0)\quad s_{t}=\alpha_{(i)}x_{i}^{l}(t)+(1-\alpha_{(i)})s_{t-1}$
This would introduce different dimensions decaying at different rates. Even with varying decay rates, because of the inherent nature of the IIR filter, we still give weightage to all values, which are never zero, unlike the FIR filter, which utilizes a window and gives no weightage to values outside the window.
Training all possible values of $\alpha$ is beyond our scope and resources. We, therefore, give the best equivalent of the EMA algorithm with our proposed method, as described in the next section.
A.2 Experiments And Results
We retain our baseline architecture precisely the same for text-8. We train for a context length of 512 with the same setup reported in our baseline section and the same dataset, with the only tweak being taking the baseline architecture and adding an EMA layer to it. We choose the number of decoder blocks to be 10, with 128 as the embedding dimension, the feed-forward dimension to be 512, and the number of heads to be 8. We opt for a two-layer feed-forward MLP inside the Transformer block after the attention block instead of a single layer typically used in Vaswani et al. (2017), with both the layers sharing the same number of neurons, i.e., 512, that of the feed-forward dimension. The final output layer of the Transformer decoder is then followed by a dense layer of 2048 neurons, followed by a dense layer of the same size as the vocabulary. This vocabulary size varies in the three modalities. For text8, it is 27, which is the number of characters plus an added extra token for space. Similar to our proposed method, we experiment with keeping half of the embedding dimensions in all the layers the same without any modifications. For the other half of the embedding dimension after all layers, we carry out EMA on 1-D signals, as described in the previous section, with $\alpha$ varying from 0 to 1 linearly for embedding dimensions 64 to 128. We see a drop in performance compared to our baseline architecture and achieve an NLL score of 0.94. For comparison, our baseline trained on text-8 scored 0.93, with our proposed method being 0.915 and 0.91 for learnable and non-learnable cases, respectively.
A.3 Discussion
There can be many reasons why EMA degrades performance. One of them can be tuning $\alpha$. There can be many possible choices, and tuning them for an expensive LLM pretraining is tough. Our proposed method, WaveletGPT, on the other hand, has a simple way of giving the weightage, which is grounded in signal processing and outperforms EMA smoothening. Further, in our learnable section, the architecture can learn the optimal weights in which, depending on the space spanned by the intermediate signals found inside LLM, it learns weights from scratch at different resolutions from the finest, i.e., window length 1 to the coarsest, i.e., window length as the context length 512.

### v4

@@ 5 Long Range Arena Benchmarks
We adapt our architecture for Long-Range Arena (LRA) tasks (Tay et al., 2021), testing on long-range prediction across text, images, and mathematical expressions. These tasks evaluate the model’s ability to handle similarity, structure, and reasoning over extended contexts. We focus on transformer-based architectures, as recently reported by (Liu et al., 2024), while other variants include state-space, hybrid models or tweaking attention mechanisms. For text, we perform binary classification on the IMDb review dataset (Maas et al., 2011) using byte-level data with a context length of 2048 to determine if a movie review is positive or negative. We use CIFAR-10 from the LRA benchmark for images, classifying sequences of 3072 pixels into one of ten categories. Lastly, we benchmark on Long ListOps, testing the model’s ability to understand hierarchically structured data in extended contexts. In our task, we use a version of ListOps of sequence lengths of up to 2K to test the ability to reason hierarchically while handling long contexts. The model needs to access all tokens and model the logical structure of the inputs to make a prediction. The task is a ten-way classification task and is considerably challenging. We use the setup provided by (Khalitov et al., 2022) to extract the data and be uniform with other benchmarks. We use identical architecture for all three modalities, only modifying the embedding matrix to accommodate different tokenizers and output categories. Our baseline consists of a 6-layer causal Transformer decoder with a model dimension of 32 and a feed-forward dimension four times that of the embedding dimension. We extract the last token of the sequence as a 32-dimensional embedding for classification, followed by a dense layer with 2048 neurons and a final dense layer corresponding to the number of categories. The input goes through an embedding layer that converts discrete tokens into a 32-dimensional vector. The input vocabularies are 256 for text/image and 16 for ListOps. The context lengths are 2048, 3072, and 1999 tokens, with 2, 10, and 10 as output categories. In our modified architecture, we introduce our waveletGPT module between each decoder layer, retaining half of the embedding dimensions as they are. For the other half, we use non-learnable kernels, increasing the kernel size from 2, 4, and 8 to 512 linearly for dimensions 16 to 32 while maintaining the causality assumption. This introduces highways that hierarchically process data at each embedding and Transformer decoder layer without adding parameters, similar to our approach for pre-trained LLM. As shown in Table 1, we achieve notable gains across all three modalities, where even minor improvements are worth reporting. We significantly outperform non-causal signal processing based methods, such as (Zhuang et al., 2024), with nearly 2% improvement on ListOps and 4.5% on a much smaller architecture: ours has 32 dimensions and six layers compared to 128 dimensions/eight layers. We limit our comparison method for fairness only with vanilla Transformer architectures. We also compare two non-casual architectures incorporating signal processing-based ideas: FNET and WavSPA. Compared to non-causal FNet, our model significantly outperformed all three LRA tasks, achieving 20% improvement on ListOps and Image and 10% on text. The most notable gain is in the ListOps task, which involves modelling a hierarchical, tree-like structure of math operations, making our model particularly suitable. To the best of our knowledge (Liu et al., 2024), this is the best result by a simple attention-based Transformer on LRA tasks.
@@ 6 Conclusion and Future Work
We showcase a powerful incorporation of a core signal processing idea, namely wavelets, into large language model pre-training. By imposing a multi-scale structure onto every intermediate embedding, we achieve the same performance 40-60% faster than a baseline architecture without adding any parameters. We achieve a substantial performance boost if we train for the same number of steps. Further we show strong gains in LRA benchmarks without adding any parameters by giving every next token prediction access to multi-scale embeddings in every decoder layer. Our method generalizes across three modalities: raw text, symbolic music, and raw audio, giving similar performance speedups on several input representations, namely raw audio samples, acoustic tokens, MIDI tokens, byte text, math expressions, BPE tokens for text, raw image pixels and characters. This shows it is generic enough for improving pre-training performance across datasets and input representations.
7 Acknowledgment
@@
This work was supported by the Stanford Institute of Human-Centered AI (Stanford-HAI) through a Google Cloud computing grant for the academic year 2023. The author is thankful to Prof. Dan Jurafsky for feedback and helpful discussions on the problem and its impact on the field.